The length of a rectangle is five inches more than its width. The perimeter is 66 inches. Find the dimensions of the rectangle.
Answer by actuary(112)

Let x equal the width of the rectangle.
Let y equal the length of the rectangle.
The description of the problem leads to the equations
l=w+5 Equation #1)
2*l+2*w=66 in. Equation #2)
Substitute equation #1 into equation #2
2*(w+5)+2*w=66 in.
Solve the equation for w
2w+10+2w=66 in.
Collect like terms
2w+2w=66-10
4w+56
Solve for w
w=56/4 = 14
So, from equation #1) l=w+5=14+5=19
I hope that this helps.